A uniform bar 15 meter long is balanced on a pivot placed at it's mid point. a boy boy of mass 55 kilograms sits on one or in the bar at a point 5 meters away from the pivot. What mass can be placed 2meter away from the other end of the bar to keep the bar horizontal


1
Expert's answer
2021-01-13T11:37:13-0500

To keep the bar horizontal, the torques on the right and left sides must be equal:


"F_1d_1=F_2d_2,\\\space\\
F_2=F_1\frac{d_1}{d_2}=(55\cdot9.8)\frac{5}{2}=1347.5\text{ N}."

Convert this force of gravity into mass:


"F_2=m_2g,\\\space\\
m_2=\frac{F_2}{g}=\frac{1347.5}{9.8}=137.5\text{ kg}."